# A Portable and Stealthy Inaudible Voice Attack Based on Acoustic Metamaterials

### v1

@@ 3       $Activate\ iwlist$
5       if $\textnormal{{feedbackMechanism(}}\textnormal{\emph{}}\textnormal{{)}}\ \textbf{==}\ true$  then
@@
10
11Function feedbackMechanism():
       // $Achieve\ feedback\ mechanism$
@@ 19       $Receive\ hotspot\ list\ L_{1}$
20       $Send\ FeedbackCmd_{1}\ and\ receive\ hotspot\ list\ L_{2}$
@@ 21       $\textbf{return}\ L_{1},L_{2}$
22
23Function Step2($L_{1},L_{2}$):
       // $If\ any\ hotspot\ signal\ disappears,\ moves\ to\ Step3$
24       $Send\ FeedbackCmd_{2}\ and\ receive\ hotspot\ list\ L_{3}$
@@ 30             $\textbf{return}\ false$
32
33
34
35Function Step3($L_{1},L_{3},dif_{1}$):
       // $If\ hotspot\ signal\ reappears,\ attack\ is\ successful$
36       $Send\ FeedbackCmd_{1}\ and\ receive\ hotspot\ list\ L_{4}$
@@ 40             Step4($Target\ list,L_{4},L_{1}$)
41             $\textbf{return}\ true$
42      else
43             $\textbf{return}\ false$
44
@@ 43             $\textbf{return}\ false$
46
47
48Function Step4($Target\ list,L_{4},L_{1}$):
       // $Restore\ the\ hotspot\ status\ of\ the\ victim\ device$
49       if $L_{4}-L_{1}==Target\ list$ then
50             $Send\ FeedbackCmd_{2}$
51
52       end if
53
54
Algorithm 1 The framework of attack overview
@@ IV-A Design Goals and Challenges
To implement a covert long-range inaudible attack system, we need to achieve two goals:
Design Goal 1. Filter out audible leakage. Maintaining the stealth nature of the attack requires filtering out audible leakage in the 100 Hz - 4000 Hz frequency range since the energy of audible leakage is concentrated in this spectrum. This filtering process is critical to maintaining the stealth nature of inaudible attacks and preventing their detection by acoustic leakage.
Design Goal 2. Increase the range of inaudible commands. To further achieve long-range attacks, we need to increase the beamforming effect of the inaudible acoustic attack command. This is because when beamforming is improved, the sound waves can be more precisely focused in the direction of a specific victim device, reducing energy dispersion and loss, which allows the sound waves to travel farther.
To achieve these two design goals, the system design must address two challenging issues.
Design Challenge 1. How to design an acoustic metamaterial with a fixed geometrical structure, which ensures that the sound signal is not leaked (for Design Goal 1) but also allows the inaudible sound signal to propagate over a longer distance (for Design Goal 2).
@@ IV-B Filter Out Audible Leakage
To address Challenge 1, we introduce a novel Leak-shielding Acoustic Metamaterial (LAM). The basic filtering mechanism of the LAM is shown in Fig. 5a. Its internal structure features a circular hole-like design (with a diameter d of 22.5 mm) surrounded by a spiral path (with a width D of 100 mm and a pitch P of 34.542 mm) [25]. When sound waves of certain frequencies pass through the metamaterial, the path variances induced by its internal structure (e.g., path 1 and path 2) result in phase opposition, leading to filtration by reflection [42, 43, 44, 45]. However, since the current structure generates a limited number of paths, it can only provide narrowband filtering (600 Hz - 1900 Hz). This limits its effectiveness in eliminating audible leakage to 100 Hz - 4000 Hz. To effectively counteract audible leakage, our LAM has undergone significant development:
@@
Expanding the filtering frequency range to 100 Hz - 4000 Hz. To further expand the LAM’s filtering range, we leverage a key observation in existing works: an opening at the connection between the circular hole-like structure and the surrounding spiral path can expand the LAM’s filtering range [25]. The reason for this observation is that an opening creates more paths, allowing more frequencies to be filtered out through phase opposition. One question that naturally arises is how to determine the opening geometric size. Through our simulation in COMSOL, we find that the height (h) and the angle ($\gamma$) of the opening both affect the filtering range of LAM. Therefore, we have tried various combinations of h and $\gamma$, then as shown in Fig. 5b, we find the opening height $h=3~{}mm$ and angle $\gamma={68.7^{\circ}}$ can expand the filtering range to 100 Hz - 4000 Hz.
Figure 5: (a) The filtering principle of LAM. (b) LAM- Leak-shielding acoustic metamaterial
Even though our LAM can utilize its unique structure to filter out audible leakage in front of the loudspeaker array through phase opposite (as shown in Fig. 5a), it cannot prevent audible leakage from leaking from the sides or back of the loudspeaker array. To address this issue, our LAM has undergone another development:
Preventing the audible leakage from leaking to other directions. As shown in Fig. 5b. Firstly, to prevent audible leakage from leaking from the sides of the loudspeaker array, we increase the thickness (${L_{1}}$) and the length (${L_{3}}$) of the wall on both sides of the LAM ($H=50mm$) to wrap around the loudspeaker array. This causes audible leakage on both sides to be reflected by the wall. Secondly, to eliminate the audible leakage on the back, we have added a cover plate at the rear of the LAM and filled the front of the cover plate with melamine sponge (${L_{2}}$) so that the audible leakage is eliminated by the thermal viscosity loss caused by the melamine sponge. Where length ${L_{3}}=17.5~{}mm$ and its value equals the thickness of the loudspeaker array.
To investigate the effects of different wall thicknesses $L_{1}$ and melamine sponge thickness $L_{2}$ on eliminating sides and back audible leakage, we simulated them in COMSOL. Since the vast majority of voice commands have the energy to 100 Hz - 4000 Hz, we set the frequency in the range of these commands. Then, we let the sound waves emitted by the sound source pass through the LAM and calculated the energy of the sound waves on the sides and back of the LAM, measured in sound pressure level (SPL). The simulation validation results of some representative $L_{1}$ and $L_{2}$ are shown in Figure 6a and Figure 6b. To ensure that the audible leakage remains inaudible to people, prior works, such as LipRead [14], have demonstrated that the “Leakage Threshold”, should be 5 dB below the “Threshold of Hearing Curve”. As a result, we derive the “Leakage Threshold” by subtracting 5 dB from the “Threshold of Hearing Curve”, as depicted by the red line in Figure 6. The results demonstrate that when the wall thickness $L_{1}$ is greater than 2 mm and the melamine sponge thickness $L_{2}$ is greater than 40 mm, audible leakage is below the “Leakage Threshold”. Therefore, on the premise of considering the filtration efficiency and compactness of LAM, we have determined the wall thickness ${L_{1}}=2~{}mm$ and the melamine sponge thickness ${L_{2}}=40~{}mm$. The total length of the LAM is ${L_{4}=H+L_{1}+L_{2}+L_{3}}=109.5~{}mm$.
@@
Figure 6: (a) Filtering effect on sides of LAM. (b) Filtering effect back of LAM. (c) Filtering effect in front of LAM.
To verify the filtering effect of the LAM, we tested its filtering performance in COMSOL. The simulation validation results are shown in Figure 6c. The result demonstrates that the audible leakage after passing through the LAM filter is below the “Leakage Threshold”, indicating the effectiveness of our LAM in mitigating audible leakage.
@@ IV-C Increase the Range of Inaudible Commands
Figure 7: Enhancement effect of inaudible commands in LAM
Figure 8: MAM (Multi-functional Acoustic Metamaterial ($H_{2}=127.5~{}mm$).
As shown in Figure 8, LAM consists mainly of a barrier and a circular piston in the center, and its structure is similar to the piston sound source [29, 30]. When the inaudible commands pass through the LAM, the circular barrier causes them to converge into the circular hole-like structure, thus enhancing them.
However, from the simulation results, we found that the surrounding spiral path inside LAM would disrupt the incident plane wave of inaudible commands, failing to form a piston sound source (the prerequisite for transforming the LAM to a piston sound source is that the incident waveform of inaudible commands is a plane wave [46, 47]).
TABLE II: The impact of various factors on the enhancement effect of inaudible commands
@@
Figure 9: The sound field of loudspeaker array
Figure 10: The results of enhancement effect.
Figure 11: The prototype of MetaAttack
Figure 12: The covert MetaAttack
To solve this issue, we let the inaudible commands pass through a LAM under different conditions and observed its waveform. Then, we find that whether the inaudible commands can form plane waves depends on three factors: (1) the number of loudspeakers; (2) the range between the loudspeaker array and circular hole-like structure; (3) carrier frequency of inaudible commands. To investigate the impact of these factors, we simulated them in COMSOL. After simulating these factors, we obtained the max SPL under different conditions in Table II.
We traverse all the possible combinations of the number of speakers, range (R), and carrier frequency to investigate their impact. Table II provides the optimal combinations of range (R) and carrier frequency under different numbers of loudspeakers. The highest gain is achieved when 12 loudspeakers are incident at 40200 Hz, positioned at a range of 18 mm (i.e., $R=18~{}mm$) from the circular hole-like structure. We also found that when using 14 loudspeakers or more, the SPL no longer increases. This is because the 12 loudspeakers perfectly cover the circular hole-like structure in the LAM, and more loudspeakers will be disturbed by other structures. Furthermore, we conducted a comparative analysis of the various layouts of the 12 speakers in the experiment referenced as Section VI-B2. This analysis led to the identification of the optimal configuration, namely, two rows and six columns, which was found to result in the greatest possible attack range.
The final structure is depicted in Fig.  8. This final structure is referred to as the Multi-functional Acoustic Metamaterial (MAM). Its length is ${H_{2}}={L_{4}}+R=127.5~{}mm$. To verify its performance, we first simulate the sound field of the loudspeaker array. The result is shown in Fig. 12. We have proven that under this structure, the waveform of the loudspeakers is a plane wave when it reaches the circular hole-like structure. We then investigate its enhancement effect in COMSOL. As shown in Fig. 12, we find that when inaudible commands meet the above conditions and pass through the MAM, the SPL increases by 11 dB, indicating that the energy of inaudible commands has been enhanced.
@@ V-A MetaAttack Components
The prototype of MetaAttack uses commercial devices and a 3D printed MAM. Fig. 12 shows our prototype system. This metamaterial serves as the basis for MetaAttack, which filters out low-frequency audible sounds and amplifies inaudible commands. The Raspberry Pi acts as a terminal for MetaAttack, facilitating the transmission of commands to attack victims and control the servo. It also detects the hotspot or Bluetooth signal from the victim device to determine the attack’s success. The amplifier amplifies the ultrasonic inaudible commands transmitted by Raspberry Pi. The servo helps MetaAttack to attack over a long distance. A loudspeaker array (12 cm $\times$ 1.4 cm) in MAM transmits inaudible commands with a center frequency of 40 kHz.
@@ V-B Stealthiness and Portability
The stealthiness and portability of inaudible attacks play a significant role in facilitating covert operations, evading detection, targeting diverse devices, and expanding the attack surface to maximize the effectiveness and reach of these malicious activities. As shown in Fig. 12, the length, width, and height of the assembled MetaAttack are 12.75 cm, 11 cm, and 18.8 cm, respectively. Its compact size allows us to carry it conveniently. Therefore, to enhance the stealth capabilities of MetaAttack, we have consolidated all attack devices into a compact travel bag. The bag can be held in hand or slung over the shoulder when performing a silent attack. With its compact size and enhanced stealth features, the MetaAttack in its travel bag configuration is a powerful tool for silent and covert operations.
@@ VI-A Experimental Setup
TABLE III: Victim devices
@@
Figure 13: Real attack scenarios: (a) S1: The attacker attacks the victim from another building at a range of 8.5 m, (b) S2: The attacker attacks the walking victim (the victim’s speed is 1 m/s and 1.5 m/s), (c) S3: The attacker sends the inaudible command: “Call 1234567890” to achieve eavesdropping purpose. (d) S4: The attacker attacks the victim in an obstacle environment.
@@
Experiment design. We conducted a range of experiments, all of which, except experiment S3, used “Send a message to Jack” as the inaudible command to achieve the purpose of sending false information. First, experiments A1 to A5 were designed to evaluate the impact of different factors on MetaAttack. Second, experiments S1 to S4 were conducted to evaluate the attack performance and the accuracy of the feedback system in real-world scenarios. As can be seen from Fig. 13, the suitable attack distances for the S1 - S4 experimental scenarios are all over 4 m, and all of them are situated in outdoor public areas or small private spaces, making them unsuitable for using large external devices. We believe that these are common scenarios, but it is difficult to fully adapt the existing methods, either the stealth requirement is met but the attack distance is not sufficient [7] [9] [36] [8], or the attacking device has a certain range but is not stealthy [14] [15]. Finally, we compared MetaAttack to prior research on inaudible attacks to highlight its superiority. Detailed descriptions of the experiments can be found in Table V.
Our experiments were approved and overseen by the University’s IRB(Institutional Review Board) to ensure compliance with data security measures and ethical guidelines for data processing.
@@
Success rate The ratio between the number of successful attacks and the total number of attacks (30 times per experiment) [10, 9].
RSA The Range at which MetaAttack can Successfully Attack (with a success rate of over 50% ) [14].
Word accuracy The ratio of the number of correctly recognized words in inaudible commands to all words [8, 14, 15].
Precision The proportion of successful attacks correctly identified by the feedback mechanism to all successful attacks identified by the feedback mechanism [53, 54, 14].
Recall The proportion of successful attacks correctly identified by the feedback mechanism to all truly successful attacks [53, 54, 14].
@@
A1 (Sec.VI-B1) Range of attack on different victim devices between MetaAttack and loudspeaker array To test the impact of range and victim devices to MetaAttack, we let MetaAttack attack the various devices (mentioned in Table III) at different ranges.
A2 (Sec.VI-B2) Attack range by different arrangements of loudspeaker arrays To find the optimal arrangement of the 12 speakers (see Sec.IV-C) for MetaAttack, we used six common settings of the speaker array to attack the victim devices.
A3 (Sec.VI-B3) Filtering of audible leakage To test the practical filtering performance of MetaAttack in inaudible attacks, we tested the audible sound leakage in front of and back of MetaAttack at frequencies between 100 Hz - 4000 Hz.
A4 (Sec.VI-B4) MetaAttack’s enhancement functionality To test the practical performance of MetaAttack’s enhancement functionality, we compared the performance of MAM, LAM and loudspeaker array.
Performance under va- rious conditions A5 (Sec.VI-B5) Noise impact To test the impact of background noise on MetaAttack, we continuously increased the noise from 55 dB to 75 dB and tested MetaAttack’s performance in this environment.
Performance in real- world scenarios S1 (Sec.VI-C1) Long-range attacks between high-rise buildings To test MetaAttack’s long-range attack performance in real-world attack scenarios, we attacked the victim devices inside another building, with the two buildings located 8.5 m apart (as shown in Fig. 13a).
S2 (Sec.VI-C2) Attacks against walking victims To test MetaAttack’s performance in attacking against walking victims, we attacked a victim who is moving on the street, and the victim’s speed is 1 m/s and 1.5 m/s (as shown in Fig. 13b).
S3 (Sec.VI-C3) Eavesdropping attacks in indoor environments To test MetaAttack’s eavesdropping performance, we let MetaAttack send “Call 123456780” to attack the victim devices to call the attacker’s number for eavesdropping (as shown in Fig. 13c).
S4 (Sec.VI-C4) Effect of passing individuals on attack range To test MetaAttack’s performance in public environments, we attacked the victim devices in a public room, and there are people along the attack path (as shown in Fig. 13d).
S5 (Sec.VI-C5) Feedback on the results of an inaudible attack To verify whether the feedback mechanism of MetaAttack is accurate, we ran it in S1 - S4 and calculated its Precision and Recall.
Performance compared to prior research Sec. VI-D Advantages of the MetaAttack in terms of concealment To verify that MetaAttack can achieve more covert attacks, we compared the size of its devices with previous attack system.

### v2

@@ 4 SendMuteCmd,AttackCmd,activateiwlistSend\ MuteCmd,\ AttackCmd,\ activate\ iwlist
7  if $\textnormal{{feedbackMechanism(}}\textnormal{\emph{}}\textnormal{{)}}\ \textbf{==}\ true\$then
@@
12
13
14Function feedbackMechanism():
15
@@ 23  $Receive\ hotspot\ list\ L_{1}$
24  $Send\ FeedbackCmd_{1}$
25  $Wait\ time\ window\ and\ receive\ hotspot\ list\ L_{2}$
@@ 26  $\textbf{return}\ L_{1},L_{2}$
27
28Function Step2($L_{1},L_{2}$):
  // Ifanyhotspotsignaldisappears,movestoStep3If\ any\ hotspot\ signal\ disappears,\ moves\ to\ Step3
29  $Send\ FeedbackCmd_{2},\ receive\ hotspot\ list\ L_{3}$
@@ 34     $\textbf{return}\ false$
36
37
38
39Function Step3($L_{1},L_{3},dif_{1}$):
  // Ifhotspotsignalreappears,attackissuccessfulIf\ hotspot\ signal\ reappears,\ attack\ is\ successful
40  $Send\ FeedbackCmd_{1}$
41  $Wait\ time\ window\ and\ receive\ hotspot\ list\ L_{4}$
@@ 45 Step4($Target\ list,L_{4},L_{1}$)
46  return true
47
48
49Function Step4($Target\ list,L_{4},L_{1}$):
  // RestorethehotspotandvolumestatusRestore\ the\ hotspot\ and\ volume\ status
50  if $L_{4}-L_{1}==Target\ list$ then
51     $Send\ FeedbackCmd_{2}\ and\ ResetCmd$
52     ClearcallandtextmessagerecordsClear\ call\ and\ text\ message\ records
53
@@ 52     ClearcallandtextmessagerecordsClear\ call\ and\ text\ message\ records
55
56
Algorithm 1 Attack overview
@@ IV-A Design Goals and Challenges
Design goal 1. Eliminate audible leakage in the 100 Hz - 4000 Hz range to maintain the stealthiness of the attack and avoid detection by acoustic leakage.
Design goal 2. Enhance beamforming to focus inaudible commands more precisely on the target device, enabling longer-range attacks.
To achieve these goals, the system must tackle two key challenges.
Design challenge 1. How to design an acoustic metamaterial, which ensures that the sound signal is not leaked (for Design Goal 1) but also allows the inaudible sound signal to propagate over a longer distance (for Design Goal 2).
@@ IV-B Filter Out Audible Leakage
To address Challenge 1, we propose a novel Leak-shielding Acoustic Metamaterial (LAM). The basic filtering mechanism of the LAM is shown in Fig. 5a. Its internal structure features a circular hole-like design (with a diameter d of 22.5 mm) surrounded by a spiral path (with a width D of 100 mm and a pitch P of 34.542 mm) [30]. When sound waves of certain frequencies pass through the metamaterial, the path variances induced by its internal structure (e.g., path 1 and path 2) result in phase opposition, leading to filtration by reflection [51, 52, 53, 54]. However, since the current structure generates a limited number of paths, it can only provide narrowband filtering (600 Hz - 1900 Hz). This limits its effectiveness in eliminating audible leakage to 100 Hz - 4000 Hz. To effectively counteract audible leakage, our LAM has undergone significant development:
@@
Expanding the filtering frequency range to 100 Hz - 4000 Hz. To further expand the LAM’s filtering range, we leverage a key observation in existing works: an opening at the connection between the circular hole-like structure and the surrounding spiral path can expand the LAM’s filtering range [30]. The reason for this observation is that an opening creates more paths, allowing more frequencies to be filtered out through phase opposition. One question that naturally arises is how to determine the opening geometric size. Through our simulations in COMSOL, we found that both the height (h) and the angle ($\gamma$) of the opening affect the filtering range of the LAM, with some key data from the simulation results shown in Fig. 6. The leakage threshold in Fig. 6 represents the sound pressure level (SPL) of audible leakage at different frequencies that can be heard by humans. According to previous studies, such as LipRead [19], the leakage threshold should be 5 dB below the threshold of hearing curve. Therefore, we derive the leakage threshold by subtracting 5 dB from the threshold of hearing curve [19].
As shown in Fig. 6a, the opening height (h) significantly affects the filtering performance in the low-frequency range. When $h=3$ mm, the sound pressure level (SPL) can be effectively suppressed below the leakage threshold. We believe this is because low-frequency sound waves, with their longer wavelengths, tend to propagate through the large-scale internal channels of the structure. In this case, the coupling strength between the central cylindrical structure (determined by h) and the rest of the system becomes critical (see Fig. 5b). If h is too small, the coupling strength is insufficient, leading to poor filtering performance. If h is too large, excessive coupling may occur, resulting in excellent filtering performance at certain frequencies, while the filtering effect at other frequencies fails to effectively suppress leakage. A height of 3 mm strikes a good balance between reflection strength and frequency coverage, ensuring stable low-frequency filtering. In addition, as shown in Fig. 6b, the aperture angle $\gamma$ is similarly crucial for high-frequency filtering. When $\gamma=68.7^{\circ}$, the SPL in this range can also be effectively suppressed below the leakage threshold. High-frequency sound waves, with shorter wavelengths, are more prone to strong multipath interference within the spiral blade structure. The angle $\gamma$ directly influences the coupling strength between this structure and the rest of the system (see Fig. 5b). Similar to h, a $\gamma$ value that is too small or too large will cause filtering imbalance, whereas $\gamma=68.7^{\circ}$ achieves stable filtering in the high-frequency range. Therefore, we find the opening height h=3mmh=3~$m$ and angle $\gamma={68.7^{\circ}}$ can expand the filtering range to 100 Hz - 4000 Hz.
Figure 5: (a) The filtering principle of LAM. (b) LAM- Leak-shielding acoustic metamaterial.
Although the LAM can filter out audible leakage in front of the loudspeaker, it still struggles to block leakage from the sides or back. To address this, we further optimized its structure:
Preventing the audible leakage from leaking to other directions. As shown in Fig. 5b, first, to prevent audible sound waves from leaking out from the sides of the speaker array, we increase the thickness ($L_{1}$}) and length ($L_{3}$}) of the walls on the sides of the LAM to encase the speaker array. This ensures that the audible leakage from the sides is reflected back by the walls. Second, to eliminate the audible leakage reflected to the back, we fill the back of the LAM with melamine sponge ($L_{2}$}), which absorbs the sound waves through its heat-viscous loss. Notably, we first confirm that the length ${L_{3}}=17.5~mm$, which is equal to the length of the speaker array.
To investigate the effects of different wall thicknesses $L_{1}$ and melamine sponge thickness $L_{2}$ on eliminating sides and back audible leakage, we performed simulations in COMSOL. The simulation results for some representative $L_{1}$ and $L_{2}$ in suppressing side and back leakage are shown in Fig. 7a and Fig. 7b. The results in Fig. 7a and Fig. 7b show that when the wall thickness $L_{1}>2~mm$ and the melamine sponge thickness $L_{2}>40~mm$, the audible leakage is below the leakage threshold. To ensure inaudibility, it is crucial to choose an appropriate thickness to enhance the compactness of the LAM. Therefore, we determined the wall thickness ${L_{1}}=2~mm$ and melamine sponge thickness ${L_{2}}=40~mm$, which balance performance and volume, with the total length of the LAM being ${L_{4}=H+L_{1}+L_{2}+L_{3}}=109.5~mm$.
Figure 6: (a) Impact of opening height $h$ on LAM front filtering. (b) Impact of opening angle $\gamma$ on LAM front filtering.
@@
Figure 7: (a) Impact of sponge thickness on LAM back filtering. (b) Impact of wall thickness on LAM sides filtering.
@@ IV-C Increase the Range of Inaudible Commands
Figure 8: Enhancement effect of inaudible commands in LAM.
Figure 9: Multi-functional Acoustic Metamaterial (MAM).
As shown in Fig. 9, LAM consists mainly of a barrier and a circular piston in the center, and its structure is similar to the piston sound source [34, 35]. When the inaudible commands pass through the LAM, the circular barrier causes them to converge into the circular hole-like structure, thus enhancing them. However, from the simulation results, we found that the surrounding spiral path inside LAM would disrupt the incident plane wave of inaudible commands, failing to form a piston sound source (the prerequisite for transforming the LAM to a piston sound source is that the incident waveform of inaudible commands is a plane wave [55, 56]).
TABLE II: The impact of various factors on the enhancement effect
@@
Figure 10: Speaker array’s sound field
Figure 11: SPL of enhancement.
Figure 12: Prototype of system
Figure 13: Covert system
To solve this issue, we let the inaudible commands pass through a LAM under different conditions and observed its waveform. Then, we find that whether the inaudible commands can form plane waves depends on three factors: (1) the number of loudspeakers; (2) the range between the loudspeaker array and circular hole-like structure; (3) carrier frequency of inaudible commands. To investigate the impact of these factors, we simulated them in COMSOL. After simulating these factors, we obtained the max SPL under different conditions in Table II. We traverse all the possible combinations of the number of speakers, range (R), and carrier frequency to investigate their impact. Table II provides the optimal combinations of range (R) and carrier frequency under different numbers of loudspeakers. The highest gain is achieved when 12 loudspeakers are incident at 40200 Hz, positioned at a range of 18 mm (i.e., R=18mmR=18~$m$) from the circular hole-like structure. We also found that when using 14 loudspeakers or more, the SPL no longer increases. This is because the 12 loudspeakers perfectly cover the circular hole-like structure in the LAM, and more loudspeakers will be disturbed by other structures. Furthermore, we conducted a comparative analysis of the various layouts of the 12 speakers in the experiment referenced as Section VI-B2. This analysis led to the identification of the optimal configuration, namely, two rows and six columns, which was found to result in the greatest possible attack range.
The final structure is depicted in Fig.  9. This final structure is referred to as the Multi-functional Acoustic Metamaterial (MAM). Its length is ${H_{2}}={L_{4}}+R=127.5~mm$. To verify its performance, we first simulate the sound field of the loudspeaker array. The result is shown in Fig. 13. We have proven that under this structure, the waveform of the loudspeakers is a plane wave when it reaches the circular hole-like structure. We then investigate its enhancement effect in COMSOL. As shown in Fig. 13, we find that when inaudible commands meet the above conditions and pass through the MAM, the SPL increases by 11 dB, indicating that the energy of inaudible commands has been enhanced.
@@ V-A MetaAttack Components
The prototype of MetaAttack uses commercial devices and a 3D-printed acoustic metamaterial (MAM), as shown in Fig. 13. This metamaterial filters out low-frequency sounds and amplifies inaudible commands. The Raspberry Pi serves as the terminal, responsible for sending attack commands, controlling the servo motor, and running the feedback mechanism. The speaker array(12 cm × 1.4 cm) in the MAM transmits commands with a center frequency of 40 kHz.
@@ V-B Stealthiness and Portability
The stealthiness and portability of inaudible attacks are crucial for covert operations and evading detection. As shown in Fig. 13, the dimensions of the assembled MetaAttack are 12.75 cm × 11 cm × 18.8 cm, making it compact and easy to carry. To enhance its stealth capabilities, we have consolidated all the attack devices into a travel bag that can be held by hand or slung over the shoulder during silent attacks. This compact and discreet design makes MetaAttack in its travel bag configuration an effective tool for silent operations.
@@ VI-A Experimental Setup
TABLE III: Target devices
@@
Figure 14: Real attack scenarios are as follows: (a) S1: The attacker launches a remote attack from a building 8.5 meters away; (b) S2: The attacker targets a walking victim outdoors; (c) S3: The attacker sends a call command to eavesdrop; (d) S4: The attacker operates in an environment with obstacles.
@@
Experiment design. We conducted a range of experiments, all of which, except experiment S3, used “Send a message to Jack” as the inaudible command to achieve the purpose of sending false information. Experiments A1 to A5 evaluated the basic performance of MetaAttack and verified the contribution of the metamaterials. The basic performance tests include attack range (A1), selection of loudspeaker array arrangement (A2), and the impact of environmental noise (A3); the contribution of the metamaterials is assessed through their ability to suppress audible leakage (A4) and the performance improvements brought by MAM and LAM (A5). Experiments S1 to S5 evaluated the attack performance and the accuracy of the feedback system. As shown in Fig. 14, the attack distances for scenarios S1 to S4 all exceed 4 meters and are located in outdoor or small private spaces, making them unsuitable for large external devices. Existing methods typically involve a trade-off between stealthiness and attack range [11, 13, 18, 12], whereas MetaAttack performs better in this regard. Detailed experimental settings can be found in Table V.
@@
Success rate Number of successful attacks / Total attempts (30 trials per device per round) [14, 13]. For Exp. A1(VI-B1).
RSA The Range at which MetaAttack can Successfully Attack (with a success rate of over 50% ) [19]. For Exps. A2(VI-B2), A4(VI-B4), A5(VI-B5), S1(VI-C1), S2(VI-C2), S4(VI-C4).
Word accuracy Proportion of correctly recognized words in inaudible commands [12, 19, 20]. For Exps. S3(VI-C3).
Precision The proportion of successful attacks correctly identified to all successful attacks identified by the feedback mechanism [62, 63, 19]. For Exp. S5(VI-C5).
Recall The proportion of successful attacks correctly identified to all truly successful attacks [62, 63, 19]. For Exp. S5(VI-C5).
@@
A1 (Sec.VI-B1) Attack range performance Attack range of MetaAttack was evaluated in comparison to a standalone speaker array.
A2 (Sec.VI-B2) Differences in speakers’ layout Six common array layouts were tested to identify the optimal configuration (Sec. IV-C).
A3 (Sec.VI-B3) Noise impact MetaAttack’s performance was tested by varying background noise from 55 to 75 dB.
A4 (Sec.VI-B4) Leakage filtering performance The audible sound leakage of MetaAttack was measured across different directions from 100-4000 Hz, alongside tests on human perception.
Basic performance evaluation A5 (Sec.VI-B5) Metamaterials’ enhancement Enhancement performance was evaluated by comparing MAM, LAM, and speaker array.
Real-world performance evaluation S1 (Sec.VI-C1) Building-to-building attack Target is attacked from 8.5 m away in a building under varying wind speeds (Fig. 14a).
S2 (Sec.VI-C2) Attack against outdoor targets The target walking outdoors was attacked, while also verifying the attack effectiveness with the device in a pocket (Fig. 14b).
S3 (Sec.VI-C3) Eavesdropping attacks MetaAttack’s eavesdropping was assessed by triggering a call (Fig. 14c).
S4 (Sec.VI-C4) Impact of passing individuals Attacks were conducted in a room with passing individuals to explore their impact (Fig. 14d).
S5 (Sec.VI-C5) Feedback mechanism performance Measured the precision and recall of the feedback mechanism in S1-S4 scenarios.
Compared to prior research Sec. VI-D MetaAttack’s concealment advantages MetaAttack’s device size and range was benchmarked against prior systems to assess its enhanced covert capabilities.